A coil of annealed copper wire has 820 turns, the average length of which is 9 inches. If the diameter of the wire is 32 miles, calculate the total resistance of the coil at 20°C.
1
Expert's answer
2020-10-13T09:30:23-0400

So the resistance is equal to "R =\frac {p\times l} {S}" , where "p" is the resistivity of the circuit, and "S" is the cross-sectional area, i.e. "p = 0. 0172\frac {Ω × mm ^ 2} {m}" , the wire length is equal to "l = N\times L" , where "N" is the number of turns, and "L" is the length of one turn, then"l = 820\times 0.23 = 188.6 m"

"S=\pi \times \frac {d^2} 4= \pi \times \frac {51499.01^2}{4}= 2 082 992 192.59 m^2" "= 2 082 992 192.59 \times 10^6 mm^2" .

"R=\frac{0.0172\times 188.6}{ 2 082 992 192.59 \times 10^6 }""=1.56 \times 10^{-15} \Omega" .